Circular No. 8309

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 F1 (NEAT) M. Hicks reports the discovery of a comet on CCD images taken with

the 1.2-m Palomar Schmidt telescope in the course of the Near-Earth Asteroid Tracking project (discovery observation below). The comet appeared soft and diffuse compared to nearby background stars, showing a coma of diameter 5′′ elongated east-west. Following posting on the NEO Confirmation Page, J. Young reported that CCD images taken by A. Grigsby and himself with the 0.6-m reflector at Table Mountain on Mar. 19.4 show a bright 4′′-diameter coma with little central condensation and a weak fuzzy area ∼ 12′′ long spanning p.a. 240^{o}-270^{o} (but no distinct tail).

2004 UT  α_{2000}  δ_{2000}  Mag.
Mar. 18.30900  12^{h}56^{m}08.^{s}70  +23^{o}10′53′′.7  18.9

The available astrometry (including prediscovery observations back to Feb. 16), the following elliptical orbital elements, and an ephemeris appear on MPEC 2004-F39.

T = 2003 Oct. 20.2305 TT  ω = 26.^{o}5748
e = 0.450051  Ω = 111.8123 }2000.0
q = 2.450159 AU  i = 18.2481
a = 4.455248 AU n^{o} = 0.1048084 P = 9.40 years

COMETS C/2003 J8, J9, K7, K8, AND K9 (SOHO) Further to IAUC 8300, K. McGleam reports measures of additional

Kreutz sungrazing comets found on SOHO C2 website images by D. Evans (C/2003 K8) and R. Kracht (remaining comets). C/2003 J8, C/2003 K7, and C/2003 K9 were also visible in C3 images. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-F29.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 J8 May 15.613  3^{h}37.^{m}6  +16^{o}14′
C/2003 J9  15.688  3 32.8  +17 31
C/2003 K7  22.321  4 11.0  +15 35
C/2003 K8  22.410  3 57.9  +18 48
C/2003 K9  23.346  4 09.8  +16 49
2004 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green